OpenStudy (anonymous):

square root of 30 times square root of 10

OpenStudy (anonymous):

its square root of 300 or 10sqrt(3)

OpenStudy (anonymous):

30sqrt 10 3.16227766 * 30 94.8683298

OpenStudy (anonymous):

what is \[\sqrt{30}x \sqrt{10}\]

OpenStudy (anonymous):

x√300

OpenStudy (anonymous):

i mean like how can i simplify these two numbers sorry

OpenStudy (anonymous):

I tried to answer this on your earlier post http://openstudy.com/users/alexg#/users/alexg/updates/4deda5534c0e8b0bc63bcbd8

OpenStudy (anonymous):

yes i know u asked the wrong question. i didnt want to ask u again because i didnt want to be a burdeen

OpenStudy (anonymous):

What is the question, the one with x, is x under the square root?

OpenStudy (anonymous):

i asked * \[\sqrt{30} \sqrt{10}\]

OpenStudy (anonymous):

Yeah click on the link I posted above.

OpenStudy (anonymous):

y do i multiply 10

OpenStudy (anonymous):

\[\sqrt{3(10)}\sqrt{10}=10\sqrt{3}\]

OpenStudy (anonymous):

oo i get it now